OpenStudy (anonymous):

At Quaker Mills High School 50% of the students have parking privileges at the school. In the same school 50% of the students participate in sports. What is the probability that a student selected at random has parking privileges at the school and participates in sports? 1/2 1/4 1 1/5

OpenStudy (mokeira):

\[\frac{ 1 }{2 }\times \frac{ 1 }{ 2 }=answer\]

OpenStudy (anonymous):

Thanks so easy!